Sustainable Growth Rate Last year Umbrellas Unlimited Corporation had an ROE of 16% and a dividend payout ratio of 40.5%. What is the sustainable growth rate?
Answer #1

Retention ratio=1-payout ratio

=(1-0.405)=0.595

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.16*0.595)/[1-(0.16*0.595)]

=10.52%(Approx).